OpenStudy (anonymous):

Answer the following questions about the function whose derivative is given below. f′(x)=(x−5)^2(x+7) a. what are the critical points of f? b. On what intervals is f increasing or decreasing? c. at what points, if any, does f assume local maximum and minimum values?

OpenStudy (anonymous):

critical points at the zeros of the derivative

OpenStudy (anonymous):

in this example the derivative changes sign at \(-7\) but not at \(5\) since it is \((x-5)^2\)

OpenStudy (dmnts):

a) you find critical points by doing this: f′(x)=0 \[(x-5)^{2}(x+7)=0\]then: x-5=0 or x+7=0 find x, then with those x find both y (x,y) will be critical points, where function changes from decreasing to increasing and the oposite

OpenStudy (anonymous):

how do you find the y's?

OpenStudy (dmnts):

y=f(x)